OpenStudy (anonymous):

How to reduce this into row echelon form? Answer would be nice as I have done it but the answer in the book is different to my answer.

OpenStudy (anonymous):

\[\left[\begin{matrix}0 & 1 &3 &2\\ 0 & 0&5&6\\1&5&1&5\end{matrix}\right]\]

OpenStudy (helder_edwin):

1st swap rows 1 and 3

OpenStudy (helder_edwin):

then swap rows 3 and 2

OpenStudy (anonymous):

the answer i get is \[\left[\begin{matrix}1 & 0&0&-5 \\ 0&1&0&2 &\\0&0&1&0\end{matrix}\right]\]

OpenStudy (helder_edwin):

but this the reduced eschelon form

OpenStudy (helder_edwin):

never mind. we'll get there

OpenStudy (helder_edwin):

by the way. IF u typed the matrix correctly, the answer your book gives u is wrong

OpenStudy (anonymous):

did you get the same result?

OpenStudy (helder_edwin):

i got \[ \large \begin{pmatrix} 1 & 0 & 0 & 59/5\\ 0 & 1 & 0 & -8/5\\ 0 & 0 & 1 & 6/5 \end{pmatrix} \]

OpenStudy (anonymous):

thats the answer in the book!! how did you get this?

OpenStudy (helder_edwin):

did u do the first two operations i told u?

OpenStudy (helder_edwin):

u should have right now 1 5 1 5 0 1 3 2 0 0 5 6

OpenStudy (anonymous):

i see what i did wrong,

OpenStudy (anonymous):

i first do the reducing then i arrange them in echelon form. Thank you @helder_edwin

OpenStudy (helder_edwin):

not necessarily. u can perform the operations in the order u see fit. as long as u don't make mistakes, of course.

OpenStudy (helder_edwin):

u r welcome